WAC Notes August 12 2020

From PREX Wiki
Jump to: navigation, search
HOW TOs for shift crew
Expert Tools
All Expert Contacts

PREX Main << Weekly Analysis Coordinator << WAC Notes CREX << Instructions

WAC Notes August 11 2020 << >> WAC Notes August 13 2020

August 12 2020

Daily Meeting BlueJeans

Production Run List - Slug 187 - IHWP IN - Flip-RIGHT - Production/Recommissioning

  • Suspicious - Run 7657 - First run after long down for hall work, injector work, and FFB recommissioning. There is clearly still some noise in the Y bpms, but not as much as yesterday. Really bad BPM jitter and noise, causing 5000ppm raw RHRS noise and inadequate regression... suspicious. Place a 12X cut at >-0.5
  • Good - Run 7658 - Excursion, cut out 60k-80k for X and E BPMs. OPS changed gang phase which fixed up bad noise and beam instability issues. There is a large excursion at the beginning, but probably benign
  • Good - Run 7660 - BPM Ys are still broad while FFB is on, but looks better now
  • Good - Run 7661 - looks good

Non-Production Runs and Bad Runs

  • Junk - Run 7659 - During alignment check with low current on Ca48


  • FFB issues at the beginning of the shift, but Gang Phase adjustments seem to get it down to a manageable level
  • There was concern over the RHRS having worse RMS regressed asymmetry
    • There is a ~30 ppm worse RHRS width, due to LHRS having 3mm more acceptance of inelastics
    • There was a false alarm of high width due to misusing 4aX BPM instead of 1X in the regressed histogram in panguin... oops!
    • Devi took an alignment run (with no dpp...) and sees everything is almost exactly as it was in March


  • Detectors are in exactly the same spots and conditions as in March - Devi's alignment check results - https://logbooks.jlab.org/entry/3825555
  • Pedestals need to be obtained in a few days and these runs reprompted